# UniChest: Conquer-and-Divide Pre-training for Multi-Source Chest X-Ray Classification

### 18 Dec 2023 v1

@@ Abstract
Vision-Language Pre-training (VLP) that utilizes the multi-modal information to promote the training efficiency and effectiveness, has achieved great success in vision recognition of natural domains and shown promise in medical imaging diagnosis for the Chest X-Rays (CXRs). However, current works mainly pay attention to the exploration on single dataset of CXRs, which locks the potential of this powerful paradigm on larger hybrid of multi-source CXRs datasets. We identify that although blending samples from the diverse sources offers the advantages to improve the model generalization, it is still challenging to maintain the consistent superiority for the task of each source due to the existing heterogeneity among sources. To handle this dilemma, we design a Conquer-and-Divide pre-training framework, termed as UniChest, aiming to make full use of the collaboration benefit of multiple sources of CXRs while reducing the negative influence of the source heterogeneity. Specially, the “Conquer" stage in UniChest encourages the model to sufficiently capture multi-source common patterns, and the “Divide" stage helps squeeze personalized patterns into different small experts (query networks). We conduct thorough experiments on many benchmarks, e.g., ChestX-ray14, CheXpert, Vindr-CXR, Shenzhen, Open-I and SIIM-ACR Pneumothorax, verifying the effectiveness of UniChest over a range of baselines, and release our codes and pre-training models at https://github.com/Elfenreigen/UniChest.
footnotetext: †: Corresponding author.
@@ 1 Introduction
Chest X-Ray (CXR) in screening chest diseases is essential to detect and control their fatal infectious impact on human lives in broad countries (Kim and Kim, 2020). To reduce the labor costs, deep learning techniques have become prevalent for machine-assisted CXR diagnosis, driving medical imaging recognition into the new era (Shin et al., 2023). Specially, with the rapid development of pre-training models, extensive studies have been conducted and shown promise in a wide range of tasks and domains (Wang et al., 2023), drawing increasing attention in medical community.
Recently, Vision-Language Pre-training techniques has significantly improved the performance of machine-aided CXR disease diagnosis (Boecking et al., 2022; Tiu et al., 2022; Wu et al., 2023; Zhang et al., 2023). Some studies even have shown the potential of surpassing the experienced radiologists in diagnosing some chest diseases (Zhang et al., 2023). Besides, in combination with medical domain-specific knowledge, these pretrained models exhibit more reasonable explanations in the lesion grounding (Wu et al., 2023; Zhang et al., 2023). Nevertheless, it is worth noting that these works for CXR VLP only consider pre-training on a single-source dataset e.g., MIMIC-CXR of about 300K samples (Johnson et al., 2019). Recalling the practice of GPT (Dale, 2021) or CLIP (Radford et al., 2021) that utilizes billions of multi-source samples, single-source VLP inherently induces drawbacks in the disease coverage and representativeness (Johnson et al., 2019), especially under real-world medical applications.
Motivated by the above limitation, we explore building a more powerful pre-training framework by leveraging multi-source CXR data. Generally, the label space union of multiple sources can help expand the coverage of the disease categories, particularly for rare diseases. Besides, samples from different sources with diverse radiation equipment, collection standards and population distributions, may complement each other (Yang et al., 2022), which helps enhance the generalization ability of pretrained models. Unfortunately, we argue that it is still very challenging to effectively utilize multi-source CXRs, as the source heterogeneity (shown in Fig. 1) also exacerbates the complexity of the CXR disease data, which could impair the holistic improvement for all source tasks during pre-training.
@@
To address this dilemma, we propose a Conquer-and-Divide pre-training framework, termed as UniChest, which maintains the merits of multi-source data collaboration and simultaneously weakens the negative effect of the source heterogeneity by a proper training isolation. Specifically, an the “Conquer" stage of UniChest, we promote the capture of multi-source common patterns at first in parts of the model, which focuses on enhancing the feature extraction abilities. Then, at the “Divide" stage of UniChest, we introduce a mixture of experts warmed up from previous stage, together with the guidance of source contrastive learning, to squeeze the source-specific patterns into different experts, which reduces the interference among sources during pre-training. With this framework, we can achieve better efficiency and effectiveness for pre-training on the large-scale multi-source CXR data. In a nutshell, the contribution of this work can be summarized as follows:
  • We are among the first attempts to explore pre-training on the multi-source CXR data and propose a Conquer-and-Divide pre-training framework to overcome the dilemma induced by the source heterogeneity in scaling up data.
@@
  • We conduct thorough experiments to show the promise of UniChest on multiple datasets, i.e., ChestX-ray14 (Wang et al., 2017), CheXpert (Irvin et al., 2019), Vindr-CXR (Nguyen et al., 2022), Shenzhen (Jaeger et al., 2014), Open-I (Demner-Fushman et al., 2016) and SIIM-ACR Pneumothorax (kag, ), achieving new state-of-the-art performance for diverse CXR diagnosis.
2 Related Works
@@ 2.1 Deep Learning for Chest X-Ray Disease Diagnosis
Considering the labor and repeatability of human experts in CXR diagnosis, it is possible to find the computer-aided solutions powered by deep learning (Van Ginneken et al., 2001; Hosny et al., 2018; Chan et al., 2020; Kieu et al., 2020). Thereby, extensive CNN-based methods for CXRs have been explored in recent years (Meedeniya et al., 2022). For instance, Khoiriyah (Khoiriyah et al., 2020) built a network comprising of three convolutional layers and three connected layers, showing remarkable performance in automatic pneumonia detection. The enhancement by transfer learning further reduced the training cost and improved the generalization performance (Yang et al., 2020; Lee et al., 2020). Specifically, several studies utilized models pre-trained on the natural domain e.g., ImageNet (Deng et al., 2009) as initialization and finetuned the last layer (Choudhuri and Paul, 2021). In addition, post-hoc techniques can be also very beneficial to enhance the stability and accuracy of the CXR diagnosis. For example, some explorations (Hashmi et al., 2020; Chouhan et al., 2020) focus on the model ensemble, i.e., directly summarizing the multiple outputs of a series of models as the final prediction, achieving the remarkable performance.
@@ 2.2 Vision-Language Pre-training in Medical Domain
Recently, Vision-Language Pre-training (VLP) models have achieved impressive success in natural domain (Bianchi et al., 2021; Jia et al., 2021; Li et al., 2021), which then drives many extensions in the medical area and improves the ability of machine-aided medical applications. ConVIRT (Zhang et al., 2022b) made the first attempt to integrate VLP into medical models, which follows the two-stream paradigm with the bidirectional contrastive learning. GLoRIA (Huang et al., 2021) explored the fine-grained information contained in the image and report, proposing a framework for learning both global and regional representations of two modalities. BioVIL (Boecking et al., 2022) focused on the representation of radiological reports and proposed a radiology-specific text encoder along with the classical VLP paradigm. CheXzero (Tiu et al., 2022) retrained a pre-trained CLIP model (Radford et al., 2021) on the CXR data and showed considerable improvement. MedKLIP (Wu et al., 2023) extracted entities from reports and converted them to the medical-specific knowledge descriptions, which enhanced the model reasoning ability. KAD (Zhang et al., 2023) built up a medical knowledge graph to fine-tune text encoder and performed the image-text contrastive learning with paired chest X-ray pairs, showing state-of-the-art capability on common benchmarks. However, all these works mentioned above are pre-trained on a single dataset from the identical source, which overlooks the non-negligible heterogeneity problem in their corresponding application on the multi-source CXR data.
@@ 3 Methodology
Figure 2: The framework of UniChest, which consists of two training stages. During the “Conquer" stage, two modality encoders first project visual and textual representations into the common space with alignment, then feed them into the first transformer query networks for prediction. The multi-source common patterns are learnt as much as possible at this stage. During the “Divide" stage, we freeze the modality encoders and squeeze the source-specific patterns via the MoE-QN module with the guidance of the enhanced supervised loss and the source contrastive learning.
@@ 3.1 Problem Formulation
Assuming that we have a training set of N samples collected from the multiple sources, $\mathcal{D}=\{(x_{i},Y_{i},l_{i},t_{i})\}_{i=1}^{N}$, where $x_{i}$ denotes the CXR image, $Y_{i}$ is the label set indicating what diseases are found for $x_{i}$, $l_{i}$ means the source identity, and $t_{i}$ denotes the report of $x_{i}$. Note that, for those samples that have no reports, we convert $Y_{i}$ to $t_{i}$ correspondingly. Our goal is to train a vision-language pre-training model on the given multi-source data $\mathcal{D}$, which can accurately diagnose the chest diseases. Specifically, in the inference phase, the model can estimate the diseases in the given set for any CXR image.
@@ 3.2 Motivation
Generally, scaling up CXR samples for pre-training should be useful to improve generalization as discussed in the previous sections. However, we should point out that the resulting multi-source samples, on one hand, enjoy the better diversity w.r.t. training samples and label space, yet on the other hand, suffer from the non-negligible source heterogeneity issue as shown in Fig. 1. Physically, the locales and time frames of source samples can be quite diverse, as ChestX-ray14 (Wang et al., 2017) contains the samples captured from 1992 to 2015 in the U.S. while VinDr-CXR (Nguyen et al., 2022) consists of CXRs in Vietnam from 2018 to 2020. Despite visually similar CXRs in human eyes, the model can actually react very differently in the face of some imperceptive factors like light circumstances and instrument qualities (Rammuni Silva and Fernando, 2022). From the t-SNE visualization in Fig. 1, we can find that the distributions of various sources exhibit the distinct disparities.
@@
To further under the influence of the source heterogeneity, we implement the straightforward pre-training on the hybrid multi-source data and compare its performance with the pre-training on single source. Specifically, we adopt the current state-of-the-art (SOTA) KAD (Zhang et al., 2023) to conduct the pre-retraining on MIMIC-CXR dataset, and compare with the pre-training on the multiple sources (termed as KAD-multi). As shown in Fig. 3, KAD-multi does not achieve the consistent improvement, and on some certain diseases, KAD-multi even significantly lags behind the vanilla KAD.
With this observation, we re-think the early VLP paradigm that is naively applied in scaling up CXR data and overlooks the source heterogeneity issue. To handle this dilemma, we actually should allow the pre-training to capture the multi-source common patterns and simultaneously can maintain the source-specific patterns, which motivates us to incorporate the philosophy of “Conquer" and “Divide" design for pre-training.
@@ 3.3.1 The Model Architecture
Given a sample $(x_{i},Y_{i},l_{i},t_{i})$, we take the ResNet-50 as the visual backbone to encode $x_{i}$ and adopt the output of the $4$-th residual block as the image representation (Wu et al., 2023), denoting as $\Phi_{\text{image}}(\cdot)$. For the report information $t_{i}$, RadGraph (Jain et al., 2021) is used to extract key entities and filter the irrelevant words. When reports containing more than one sentence, we extract entities sentence by sentence, and use [SEP] token as the separation between different entities [^1] For CXR samples without reports, we use their corresponding labels as the input entity and also take [SEP] token to separate the multiple labels.. After the entity extraction, we use PubMedBERT (Gu et al., 2021) pre-trained on Unified Medical Language System (UMLS) data (Bodenreider, 2004) as textual backbone $\Phi_{\text{text}}$ for generating text representation (Zhang et al., 2023). For clarity, we summarize the modality-specific encoding process with equations as below,
@@
For the disease prediction, we introduce a MoE-QN Module, namely, mixture of query networks, where each query network consists of a few transformer decoder layers. The MoE-QN module plays the role of overcoming the source heterogeneity issues by squeezing the source-specific patterns into different query networks, but its training should carefully follow our Conquer-and-Divide schedules, which will be discussed in the subsequent sections. For each query network, we take the fine-grained visual representation $\boldsymbol{I}_{i}$ as Key and Value, and utilize the textual representation of the disease set $Y_{i}=\{Y_{i1},...,Y_{ic}\}$ (similarly encoded by $\Phi_{\text{text}}(\cdot)$) as Query. The output from the sequential transformer decoder will be fed to one MLP layer to obtain the prediction of each Query Network,
$s_{i}^{k}=\Phi_{\text{QN}}^{k}(\boldsymbol{I}_{i},\boldsymbol{I}_{i},\Phi_{\text{text}}(Y_{i}))\in\mathbb{R}^{c},~{}k=1,...,K,$ (2)
where $c$ is the total class number of all sources and $K$ is the number of query networks. Then, we transform $\{s_{i}^{k}\}_{k=1}^{K}$ of all query networks to the total prediction by an automatic linear combination, which will be described at the “Divide" stage.
3.3.2 The Training at the “Conquer" Stage
As discussed in the earlier sections, the source heterogeneity issue can be the core bottleneck to deteriorate the performance of pre-training for the consistent improvement for each source task. Therefore, at this “Conquer" stage, we first encourage the model to capture multi-source common patterns as many as possible. In brief, the training loss of this stage involves two parts, the image-text bidirectional contrastive loss $\mathcal{L}_{\text{BCL}}$ and the multi-label cross-entropy loss $\mathcal{L}_{\text{BCE}}$ w.r.t. the first query network prediction. Taking a batch of $M$ samples as an example, $\mathcal{L}_{\text{BCL}}$ emphasizes the alignment between the global visual representation and the textual representation, which could be formulated as follows,

### 22 Mar 2024 v2

@@ Abstract
Vision-Language Pre-training (VLP) that utilizes the multi-modal information to promote the training efficiency and effectiveness, has achieved great success in vision recognition of natural domains and shown promise in medical imaging diagnosis for the Chest X-Rays (CXRs). However, current works mainly pay attention to the exploration on single dataset of CXRs, which locks the potential of this powerful paradigm on larger hybrid of multi-source CXRs datasets. We identify that although blending samples from the diverse sources offers the advantages to improve the model generalization, it is still challenging to maintain the consistent superiority for the task of each source due to the existing heterogeneity among sources. To handle this dilemma, we design a Conquer-and-Divide pre-training framework, termed as UniChest, aiming to make full use of the collaboration benefit of multiple sources of CXRs while reducing the negative influence of the source heterogeneity. Specially, the “Conquer” stage in UniChest encourages the model to sufficiently capture multi-source common patterns, and the “Divide” stage helps squeeze personalized patterns into different small experts (query networks). We conduct thorough experiments on many benchmarks, e.g., ChestX-ray14, CheXpert, Vindr-CXR, Shenzhen, Open-I and SIIM-ACR Pneumothorax, verifying the effectiveness of UniChest over a range of baselines, and release our codes and pre-training models at https://github.com/Elfenreigen/UniChest.
{IEEEkeywords}
Chest X-Rays, Medical Imaging Diagnosis, Conquer and Divide, Vision-Language Pre-training.
@@ 1 Introduction
[…]
Chest X-Ray (CXR) in screening chest diseases is essential to detect and control their fatal infectious impact on human lives in broad countries [1]. To reduce the labor costs, deep learning techniques have become prevalent for machine-assisted CXR diagnosis, driving medical imaging recognition into the new era [2]. Specially, with the rapid development of pre-training models, extensive studies have been conducted and shown promise in a wide range of tasks and domains [3], drawing increasing attention in medical community.
Recently, Vision-Language Pre-training techniques has significantly improved the performance of machine-aided CXR disease diagnosis [4, 5, 6, 7]. Some studies even have shown the potential of surpassing the experienced radiologists in diagnosing some chest diseases [7]. Besides, in combination with medical domain-specific knowledge, these pretrained models exhibit more reasonable explanations in the lesion grounding [6, 7]. Nevertheless, it is worth noting that these works for CXR VLP only consider pre-training on a single-source dataset e.g., MIMIC-CXR of about 300K samples [8]. Recalling the practice of GPT [9] or CLIP [10] that utilizes billions of multi-source samples, single-source VLP inherently induces drawbacks in the disease coverage and representativeness [8], especially under real-world medical applications.
Motivated by the above limitation, we explore building a more powerful pre-training framework by leveraging multi-source CXR data. Generally, the label space union of multiple sources can help expand the coverage of the disease categories, particularly for rare diseases. Besides, samples from different sources with diverse radiation equipment, collection standards and population distributions, may complement each other [11], which helps enhance the generalization ability of pretrained models. Unfortunately, we argue that it is still very challenging to effectively utilize multi-source CXRs, as the source heterogeneity (shown in Fig. 1) also exacerbates the complexity of the CXR disease data, which could impair the holistic improvement for all source tasks during pre-training.
@@
To address this dilemma, we propose a Conquer-and-Divide pre-training framework, termed as UniChest, which maintains the merits of multi-source data collaboration and simultaneously weakens the negative effect of the source heterogeneity by a proper training isolation. Specifically, at the “Conquer” stage of UniChest, we promote the capture of multi-source common patterns at first in parts of the model, which focuses on enhancing the feature extraction abilities. Then, at the “Divide” stage of UniChest, we introduce a mixture of experts warmed up from previous stage, together with the guidance of source contrastive learning, to squeeze the source-specific patterns into different experts, which reduces the interference among sources during pre-training. With this framework, we can achieve better efficiency and effectiveness for pre-training on the large-scale multi-source CXR data. In a nutshell, the contribution of this work can be summarized as follows:
  • We explore pre-training on large-scale multi-source CXR data and propose a Conquer-and-Divide pre-training framework to overcome the dilemma induced by the source heterogeneity in scaling up the data.
@@
  • We conduct thorough experiments to show the promise of UniChest on multiple datasets, i.e., ChestX-ray14 [12], CheXpert [13], Vindr-CXR [14], Shenzhen [15], Open-I [16] and SIIM-ACR Pneumothorax [17], achieving new state-of-the-art performance for diverse CXR diagnosis.
2 Related Work
@@ 2.1 Deep Learning for Chest X-Ray Disease Diagnosis
Considering the labor and repeatability of human experts in CXR diagnosis, it is possible to find the computer-aided solutions powered by deep learning [18, 19, 20, 21]. Thereby, extensive CNN-based methods for CXRs have been explored in recent years [22]. For instance, Khoiriyah [23] built a network comprising of three convolutional layers and three connected layers, showing remarkable performance in automatic pneumonia detection. The enhancement by transfer learning further reduced the training cost and improved the generalization performance [24, 25]. Specifically, several studies utilized models pre-trained on the natural domain e.g., ImageNet [26] as initialization and finetuned the last layer [27]. In addition, post-hoc techniques can be also very beneficial to enhance the stability and accuracy of the CXR diagnosis. For example, some explorations [28, 29] focus on the model ensemble, i.e., directly summarizing the multiple outputs of a series of models as the final prediction, achieving the remarkable performance.
@@ 2.2 Vision-Language Pre-training in Medical Domain
Recently, Vision-Language Pre-training (VLP) models have achieved impressive success in natural domain [30, 31, 32], which then drives many extensions in the medical area and improves the ability of machine-aided medical applications. ConVIRT [33] made the first attempt to integrate VLP into medical models, which follows the two-stream paradigm with the bidirectional contrastive learning. GLoRIA [34] explored the fine-grained information contained in the image and report, proposing a framework for learning both global and regional representations of two modalities. MedCLIP [35] proposed one decoupled multimodal contrastive learning framework based on CLIP to scale the usable training data from two distinct sources. BioVIL [4] focused on the representation of radiological reports and proposed a radiology-specific text encoder along with the classical VLP paradigm. CheXzero [5] retrained a pre-trained CLIP model [10] on the CXR data and showed considerable improvement. MedKLIP [6] extracted entities from reports and converted them to the medical-specific knowledge descriptions, which enhanced the model reasoning ability. KAD [7] built up a medical knowledge graph to fine-tune text encoder and performed the image-text contrastive learning with paired chest X-ray pairs, showing state-of-the-art capability on common benchmarks. However, all these works mentioned above are either pre-trained on a single dataset from the identical source or on dual sources like MedCLIP [35], which overlook the non-negligible heterogeneity problem in their corresponding application on the multi-source CXR data.
@@ 3.1 Problem Formulation
Assuming that we have a training set of N samples collected from the multiple sources, $\mathcal{D}=\{(x_{i},Y_{i},l_{i},t_{i})\}_{i=1}^{N}$, where $x_{i}$ denotes the CXR image, $Y_{i}$ is the label set indicating what diseases are found for $x_{i}$, $l_{i}$ means the source identity, and $t_{i}$ denotes the report of $x_{i}$. Note that, for those samples that have no reports, we convert $Y_{i}$ to $t_{i}$ correspondingly. Our goal is to train a vision-language pre-training model on the given multi-source data $\mathcal{D}$, which can accurately diagnose the chest diseases. Specifically, in the inference phase, the model can estimate the diseases in the given set for any CXR image. Here, we would like to clarify that our method adheres to the vision-language pre-training paradigm instead of transition supervised learning paradigm, since it inherently extracts representation vectors from images and text data for contrastive pre-training and can perform to open-set and zero-shot evaluations. Conventional supervised learning actually cannot perform such evaluations due to the rigid dimension issue about the prediction as well as the generalization dilemma towards the semantics of new classes.
@@ 3.2 Motivation
Generally, scaling up CXR samples for pre-training should be useful to improve generalization as discussed in the previous sections. However, we should point out that the resulting multi-source samples, on one hand, enjoy the better diversity w.r.t. training samples and label space, yet on the other hand, suffer from the non-negligible source heterogeneity issue as shown in Fig. 1. Physically, the locales and time frames of source samples can be quite diverse, as ChestX-ray14 [12] contains the samples captured from 1992 to 2015 in the U.S. while VinDr-CXR [14] consists of CXRs in Vietnam from 2018 to 2020. Even though CXRs might appear indistinguishable to the human eye, the diagnostic models can actually respond very differently in the face of some imperceptive factors like the dosage of X-ray used and the quality of the imaging instrument [36]. From the t-SNE visualization in Fig. 1, we can find that the distributions of various sources exhibit the distinct disparities.
@@
To further under the influence of the source heterogeneity, we implement the straightforward pre-training on the hybrid multi-source data and compare its performance with the pre-training on single source. Specifically, we adopt the current state-of-the-art (SOTA) KAD [7] to conduct the pre-retraining on MIMIC-CXR dataset, and compare with the pre-training on the multiple sources (termed as KAD-multi). As shown in Fig. 2, KAD-multi does not achieve the consistent improvement, and on some certain diseases, KAD-multi even significantly lags behind the vanilla KAD.
With this observation, we re-think the early VLP paradigm that is naively applied in scaling up CXR data and overlooks the source heterogeneity issue. To handle this dilemma, we actually should allow the pre-training to capture the multi-source common patterns and simultaneously can maintain the source-specific patterns, which motivates us to incorporate the philosophy of “Conquer” and “Divide” design for pre-training.
@@ 3.3.1 The Model Architecture
Given a sample $(x_{i},Y_{i},l_{i},t_{i})$, we take the ResNet-50 as the visual backbone to encode $x_{i}$ and adopt the output of the $4$-th residual block as the image representation [6], denoting as $\Phi_{\text{image}}(\cdot)$. For the report information $t_{i}$, RadGraph [37] is used to extract key entities and filter the irrelevant words. When reports containing more than one sentence, we extract entities sentence by sentence, and use [SEP] token as the separation between different entities [^1] For CXR samples without reports, we use their corresponding labels as the input entity and also take [SEP] token to separate the multiple labels.. After the entity extraction, we use PubMedBERT [38] pre-trained on Unified Medical Language System (UMLS) data [39] as textual encoder backbone $\Phi_{\text{text}}$ for generating text representation [7]. For clarity, we summarize the modality-specific encoding process as below,
@@
For the disease prediction, we introduce a MoE-QN Module, namely, mixture of query networks, where each query network consists of a few transformer decoder layers (4-layers in default). The MoE-QN module plays the role of overcoming the source heterogeneity issues by squeezing the source-specific patterns into different query networks, but its training should carefully follow our Conquer-and-Divide schedules, which will be discussed in the subsequent sections. For each query network, we take the fine-grained visual representation $\boldsymbol{I}_{i}$ as Key and Value, and utilize the textual representation of the disease set $Y_{i}=\{Y_{i1},...,Y_{ic}\}$ (encoded by $\Phi_{\text{text}}(\cdot)$ similarly as in Eq. (1)) as Query. The output from the sequential transformer decoder will be fed to one MLP layer to obtain the prediction of each Query Network,
${s_{i}^{k}=\Phi_{\text{QN}}^{k}(\boldsymbol{I}_{i},\Phi_{\text{text}}(Y_{i}))\in\mathbb{R}^{c},~{}k=1,...,K,}$ (2)
where $c$ is the total class number of all sources and $K$ is the number of query networks. Then, we transform $\{s_{i}^{k}\}_{k=1}^{K}$ of all query networks to the total prediction by an automatic linear combination, which will be described at the “Divide” stage.
Figure 3: The framework of UniChest, which consists of two training stages. During the “Conquer” stage, two modality encoders first project visual and textual representations into the common space with alignment, then feed them into the first transformer query networks for prediction. The multi-source common patterns are learnt as much as possible at this stage. During the “Divide” stage, we freeze the modality encoders and squeeze the source-specific patterns via the MoE-QN module with the guidance of the enhanced supervised loss and the source contrastive learning.
3.3.2 The Training at the “Conquer” Stage
As discussed in the earlier sections, the source heterogeneity issue can be the core bottleneck to deteriorate the performance of pre-training for the consistent improvement for each source task. Therefore, at this “Conquer” stage, we first encourage the model to capture multi-source common patterns as many as possible following the prevalent vision-language pre-training framework like KAD [7]. The training loss of this stage involves two parts, the image-text bidirectional contrastive loss $\mathcal{L}_{\text{BCL}}$ and the specially enhanced multi-label cross-entropy loss $\mathcal{L}_{\text{BCE}}$ (see Section IV.B for its implementation detail) w.r.t. the first query network prediction. Taking a batch of $M$ samples as an example, $\mathcal{L}_{\text{BCL}}$ emphasizes the alignment between the global visual representation and the textual representation, which could be formulated as follows,